OpenStudy (anonymous):

Find the Volume of a right cone with radius 3m and height 8m. Leave answer in terms of pi please.

OpenStudy (anonymous):

substitute it with the volume of a cone]

OpenStudy (anonymous):

The volume of a cone is = (1/3)*pi*r^2*h =(1/3)*pi*(3^2)(8) =24pi

OpenStudy (anonymous):

1/3 *pi *9*8=24pi

Directrix (directrix):

@Dido525 OpenStudy values the Learning process - not the ‘Give you an answer’ process •Don’t post only answers - guide the asker to a solution. http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

ok got it sir